OpenStudy (anonymous):

Find the 10th term of the geometric sequence. -0.125, 0.25, -0.5, 1, -2, ...

OpenStudy (solomonzelman):

\(\large\color{black}{ a_n=a_1\times r^{n-1} }\) is the geometirc formula for a(n), so it would be logical to say that: \(\large\color{black}{ a_{10}=a_1\times r^{10-1} }\) which gives us: \(\Large\color{red}{ a_{10}=a_1\times r^{9} }\)

OpenStudy (solomonzelman):

you just need to put in the common ratio and the first term into the red equation.

OpenStudy (anonymous):

i keep getting the wrong answer :(

OpenStudy (anonymous):

i thought the ratio was -.5 because a1/a2?

OpenStudy (solomonzelman):

yes, yes. my bad.

OpenStudy (solomonzelman):

it is just -2.

OpenStudy (solomonzelman):

I was so off. Sorry for confusing you. Saturday must be a very bad day....

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_{10}=a_{1}\times r^9 }\) \(\large\color{blue}{ a_{10}=a_{1}\times 2^9 }\)

OpenStudy (solomonzelman):

you have the a(1) left.

OpenStudy (anonymous):

im confused!

OpenStudy (anonymous):

why is it just -2?

OpenStudy (xapproachesinfinity):

well the ratio of a geometric sequence is the number you multiply previous to get the next in your case the first term is -0.125 to get to 0.25 we multiplied by -2 then again we multiply 0.25 by -2 to get the third term -0.5 and so on

OpenStudy (xapproachesinfinity):

the formula that generate all the terms is given by \(A_n=A_1r^{n-1}\) like @SolomonZelman mentioned above to get the tenth term just evaluate \(A_{10}=-0.125*(-2)^9\)

OpenStudy (anonymous):

and its -64

OpenStudy (anonymous):

but i am still stuck on the ratio :(

OpenStudy (xapproachesinfinity):

- i don't think so, it shouldn't be negative

OpenStudy (anonymous):

why not?

OpenStudy (xapproachesinfinity):

to get the ratio of a geometric sequence all you need is divide the the next term by the previous term 0.25/-0.125 which is -2

OpenStudy (anonymous):

but i thought it was a1/a2

OpenStudy (xapproachesinfinity):

it shouldn't be negative since the power is odd \((-2)^9\) this should be negative and the with -0.125 the negative sign should be gone

OpenStudy (xapproachesinfinity):

hmm no! try -.5 it won't give you 0.25 there is no succession in term

OpenStudy (anonymous):

okay thank you

OpenStudy (xapproachesinfinity):

i did get 64 but not negative

OpenStudy (xapproachesinfinity):

welcome^_^ where did you learn that it is a1/a2?

OpenStudy (anonymous):

a video in my course :(

OpenStudy (xapproachesinfinity):

hmmm probably doing a different problem but the successive of terms is kept

OpenStudy (anonymous):

probably ima look at it again.

OpenStudy (anonymous):

but thank you so much for explaining it

OpenStudy (xapproachesinfinity):

welcome!

OpenStudy (xapproachesinfinity):

hmm i think they are doing a1/a2 to find 1/r not r and then flip 1/r to get the right ratio lol i guess or the problem is set differently all you need to keep in mind is that the ratio is what need to multiply the previous to get next term

OpenStudy (xapproachesinfinity):

good luck pal:)

OpenStudy (anonymous):

okay thank you (: